Puget Sound Divers is a company that provides diving services such as underwater ship repairs to clients in the Puget Sound area. The company’s planning budget for May appears below:

Puget Sound Divers
Planning Budget
For the Month Ended May 31
Budgeted diving-hours (q) 100
Revenue ($365.00q) $ 36,500
Expenses:
Wages and salaries ($8,000 + $125.00q) 20,500
Supplies ($3.00q) 300
Equipment rental ($1,800 + $32.00q) 5,000
Insurance ($3,400) 3,400
Miscellaneous ($630 + $1.80q) 810
Total expense 30,010
Net operating income $ 6,490

During May, the company’s actual activity was 105 diving-hours.

Required:

Prepare a flexible budget for May.
Answer #1

105 driving hours $38,325.00 Actual activity Revenue Less: Expenses Wages and salaries Supplies Equipment rent Insurance Misc

Working note:

Actual activity 105 driving hours Revenue =105*365 Less: Expenses Wages and salaries =8000+(105*125) Supplies =105*3 Equipmen